OpenStudy (anonymous):

32^t+1=16^1-t cant seem to get the right solution

OpenStudy (anonymous):

is 16 to the power of 1-t ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I use a common base and then distribute

OpenStudy (anonymous):

32 = 2^5 16 = 2^4

OpenStudy (anonymous):

that is, solve \[5(t+1)=4(1-t)\] for t